A kit has an 8 inch side and a 15 inch side which form a right angle. find the length of the diagonals of the kite.: A kit has an 8 inch side and a 15 inch side which form a right angle. find the length of the diagonals of the kite.
Answer by nerdybill(1129)

A kite has an 8 inch side and a 15 inch side which form a right angle. find the length of the diagonals of the kite.
.
If a triangle has a "right angle", then you can apply Pythagorean theorem:
Let x = diagonal of kite
then
x^2 = 8^2 + 15^2
.
Now, solving for 'x':
x^2 = 8^2 + 15^2
x^2 = 64 + 225
x^2 = 289
x = 17 inches